OpenStudy (anonymous):

solve for the variable in the following equation 2/3x + 5/6x =1/2x+10

OpenStudy (anonymous):

cmon helppppp

OpenStudy (anonymous):

\[\frac{2}{3}x+\frac{5}{6}x=\frac{1}{2}x+10\]

OpenStudy (anonymous):

first multiply by somethin to get rid of the fractions

OpenStudy (anonymous):

in order to get rid of the highest denominator you'd have to multiply by _